OpenStudy (aglna):

Can someone help me Convert 67,240,00 into scientific notation

OpenStudy (mathstudent55):

In scientific notation, a number is represented by a number from 1 to under 10 multiplied by a integer power of 10. Examples of numbers in scientific notation: \(3.45 \times 10^3\) \(1 \times 10^{-4} \) \(9.98 \times 10^{-10} \) You need a number from 1 to under 10. Use the digits of the number you are given, and place the decimal point in a way that you end up with a number from 1 to less than 10. Don;t worry about the power of ten yet. Can you do that?

OpenStudy (cas_fangirl_14):

6.724x10^7 Like instead of writing 0.0000000056, we write 5.6 x 10^-9

OpenStudy (mathstudent55):

BTW, are you missing a zero at the end of your number, or are the commas in the wrong positions?

OpenStudy (aglna):

Yes

OpenStudy (mathstudent55):

Yes, which? 67,240,00 should be 67,240,000 ?

OpenStudy (cas_fangirl_14):

Then the answer would be 6.724x10^8

OpenStudy (mathstudent55):